OpenStudy (anonymous):

f(x)=1-6x Solve f(x)=0

OpenStudy (saifoo.khan):

f(x)=1-6x As it says that the function is equal to zero, so: 0 = 1 - 6x Can you solve it now?

OpenStudy (anonymous):

is it x= 1/6?

OpenStudy (saifoo.khan):

Right! :)

OpenStudy (anonymous):

how would you show working out

OpenStudy (katielong):

0=1-6x -1=-6x -1/-6=x 1/6=x lol

OpenStudy (katielong):

just write that down